# Exponentially accelerated relaxation and quantum Mpemba effect in open quantum systems

### v1

@@ IV.2 Quantum relative entropy
The right-hand side of Eq. (25) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{g_{l}}$ of the nonnegative elements ${g_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\varepsilon)$. To validate the inequality in Eq. (25), we apply the same reasoning as in Sec. IV.2. The ideia is that there exist an optimal permutation matrix $P_{\pi,\text{opt}}$} that satisfies the lower bound $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\geq{g_{\text{max}}}\geq{\sum_{l}}\,{\epsilon_{l}}{g_{l}}$, where ${g_{\text{max}}}=\text{max}\{{g_{1}},{g_{2}},\ldots,{g_{d}}\}$. We note that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)={\sum_{l}}\,{\alpha_{l}}{\lambda_{\pi(l)}}$, where $\pi(\bullet)$ maps index $l$ to $\pi(l)$, with $l,\pi(l)\in\{1,2,\ldots,d\}$. Therefore, to ensure that ${\sum_{l}}\,{\alpha_{l}}{\lambda_{\pi(l)}}\geq{g_{\text{max}}}\geq{\sum_{l}}\,{\epsilon_{l}}{g_{l}}$, it is sufficient to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\geq{\lambda_{\pi(2)}}\geq\ldots\geq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{n}}$ are listed in ascending order, or vice-versa Hardy et al. (1988). This proves the inequality in Eq. (22), which means that there exists $P_{\pi}$ that maximizes QRE evaluated for ${\rho^{\prime}}({t_{0}})$ and $\texttt{R}_{1}$}.
@@ IV.3 Trace distance
with $|A|=\sqrt{{A^{\dagger}}A}$. Overall, TD is (i) non-negative, i.e., ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)={\mathcal{D}_{\text{TD}}}(\varrho,\rho)$; (iii) unitarily invariant, ${\mathcal{D}_{\text{TD}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{TD}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator; (iv) monotonically decreasing under completely positive and trace preserving (CPTP) maps, i.e., ${\mathcal{D}_{\text{TD}}}\left(\mathcal{E}(\rho),\mathcal{E}(\varrho)\right)\leq{\mathcal{D}_{\text{TD}}}(\rho,\varrho)$, with $\mathcal{E}(\bullet)$ being a given CPTP operation Nielsen and Chuang (2010).
Figure 2: (Color online) The quantum Mpemba effect in the dynamics of the transverse field Ising (TFI) model with open boundary conditions [see details in Sec. V], where $N=5$ spins, $h=J/2$, $\gamma=1$, and $T=0.1$. To investigate the relaxation of the instantaneous state $\rho(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) Hilbert-Schmidt distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$; (c) trace distance, ${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $Jt$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
@@
where ${u^{\uparrow}}=\text{diag}({\lambda_{1}},\ldots,{\lambda_{d}})$ defines a vector formed by the eigenvalues of $\rho({t_{0}})$ listed in ascending order, while ${v^{\downarrow}}=\text{diag}({q_{1}},\ldots,{q_{d}})$ is a vector that contains the eigenvalues of $\texttt{R}_{1}$ listed in descending order, with ${q_{l}}=e^{-{\varepsilon_{l}}/{k_{B}}T}/Z$ being the $l$-th eigenvalue of the steady state. Next, note that the trace distance for the initial dressed state ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$ is written as
@@
where we used the fact the TD is unitarily invariant. Here, $\Sigma=\text{diag}({q_{1}},\ldots,{q_{d}})$, while ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots{\lambda_{\pi(d)}})$ is the matrix obtained by rearranging the entries of the diagonal matrix $D$ that contains the eigenvalues of the initial state. It is worth noting that always exist an appropriate permutation matrix $P_{\pi}$} that rearranges the eigenvalues of the initial state such that ${\mathcal{D}_{\text{TD}}}({P_{\pi}}{D}{P_{\pi}^{\dagger}},\Sigma)={\mathcal{D}_{\text{TD}}}({u^{\uparrow}},{v^{\downarrow}})$. In this case, combining Eqs. (27) and (IV.3), it yields

### v2

@@ IV.2 Quantum relative entropy
To validate the inequality in Eq. (25), we apply the same reasoning as in Sec. IV.2. The right-hand side of Eq. (25) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{g_{l}}$ of the nonnegative elements ${g_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\varepsilon)$. The ideia is that there exist an optimal permutation matrix $P_{\pi,\text{opt}}$} that satisfies the lower bound $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\varepsilon)\geq{g_{\text{max}}}\geq{\sum_{l}}\,{\epsilon_{l}}{g_{l}}$, where ${g_{\text{max}}}=\text{max}\{{g_{1}},\ldots,{g_{d}}\}$. We note that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\varepsilon)={\sum_{l}}\,{\varepsilon_{l}}{\lambda_{\pi(l)}}$. Therefore, to ensure that ${\sum_{l}}\,{\varepsilon_{l}}{\lambda_{\pi(l)}}\geq{g_{\text{max}}}\geq{\sum_{l}}\,{\epsilon_{l}}{g_{l}}$, it is sufficient to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\geq\ldots\geq{\lambda_{\pi(d)}}$ being listed in descending order, provided that ${\varepsilon_{1}}\geq\ldots\geq{\varepsilon_{n}}$ are listed in descending order, or vice-versa Hardy et al. (1988). This proves the inequality in Eq. (22), which means that there exists $P_{\pi}$ that maximizes QRE evaluated for ${\rho^{\prime}}({t_{0}})$ and $\texttt{R}_{1}$}.
@@ IV.3 Trace distance
with $|X|:=\sqrt{{X^{\dagger}}X}$. Overall, TD is (i) non-negative, i.e., ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)\geq 0$, with ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${\mathcal{D}_{\text{TD}}}(\rho,\varrho)={\mathcal{D}_{\text{TD}}}(\varrho,\rho)$; (iii) unitarily invariant, ${\mathcal{D}_{\text{TD}}}(V\rho{V^{\dagger}},V\varrho{V^{\dagger}})={\mathcal{D}_{\text{TD}}}(\rho,\varrho)$, where ${V^{\dagger}}={V^{-1}}$ is a unitary operator; (iv) monotonically decreasing under completely positive and trace preserving (CPTP) maps, i.e., ${\mathcal{D}_{\text{TD}}}\left(\mathcal{E}(\rho),\mathcal{E}(\varrho)\right)\leq{\mathcal{D}_{\text{TD}}}(\rho,\varrho)$, with $\mathcal{E}(\bullet)$ being a given CPTP operation Nielsen and Chuang (2010).
@@
where ${u^{\uparrow}}=\text{diag}({\lambda_{1}},\ldots,{\lambda_{d}})$ defines a matrix formed by the eigenvalues of $\rho({t_{0}})$ listed in ascending order, while ${v^{\downarrow}}=\text{diag}({\alpha_{1}},\ldots,{\alpha_{d}})$ is a matrix that contains the eigenvalues of $\texttt{R}_{1}$ listed in descending order, with ${\alpha_{l}}=e^{-{\varepsilon_{l}}/{k_{B}}T}/Z$ being the $l$-th eigenvalue of the steady state. Next, note that the trace distance for the initial dressed state ${\rho^{\prime}}({t_{0}})$ and the steady state $\texttt{R}_{1}$ is written as
@@
where we used the fact the TD is unitarily invariant. Here, $\Sigma=\text{diag}({\alpha_{1}},\ldots,{\alpha_{d}})$, while ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots{\lambda_{\pi(d)}})$ is the matrix obtained by rearranging the entries of the diagonal matrix $D$ that contains the eigenvalues of the initial state. It is worth noting that always exist an appropriate permutation matrix $P_{\pi}$} that rearranges the eigenvalues of the initial state such that ${\mathcal{D}_{\text{TD}}}({P_{\pi}}{D}{P_{\pi}^{\dagger}},\Sigma)={\mathcal{D}_{\text{TD}}}({u^{\uparrow}},{v^{\downarrow}})$. In this case, combining Eqs. (27) and (IV.3), it yields
@@ V Examples
In this section, we illustrate our findings for the case of the dynamics of (i) two-level systems [see Sec. V.1]; and (ii) many-body quantum systems [see Sec. V.2], particularly addressing the transverse field Ising model, and also the XXZ model.
V.1 Two-level system
We consider a two-level system ($d=2$) weakly coupled to a bosonic reservoir at temperature $T$. The Hamiltonian of the system is given by $H={U_{1}}{\varepsilon}{U_{1}^{\dagger}}$, where ${\varepsilon}=\text{diag}({\varepsilon_{1}},{\varepsilon_{2}})$, with ${\varepsilon_{1}}>{\varepsilon_{2}}$, while ${U_{1}}=i(|{1}\rangle\langle{0}|-|{0}\rangle\langle{1}|)$. The vectors $\{|{0}\rangle,|{1}\rangle\}$ are the eigenstates of Pauli matrix $\sigma_{z}$}, i.e., ${\sigma_{z}}|{x}\rangle={(-1)^{x}}|{x}\rangle$, for $x=\{0,1\}$, and define the computational basis. The system is initialized at ${t_{0}}=0$ in the pure single-qubit state $\rho({t_{0}})={\Lambda}{D}{\Lambda^{\dagger}}$, where $D=\text{diag}(0,1)=|1\rangle\langle{1}|$, and $\Lambda={-i}|{0}\rangle\langle{-}|+|{1}\rangle\langle{+}|$, with $|{\pm}\rangle=(1/\sqrt{2})(|{0}\rangle\pm|{1}\rangle)$. The effective dynamics of the two-level system is governed by the Liouvillian $\mathcal{L}[\bullet]$ in Eq. (1) [see also Eq. (2)], with in turn is related to the Lindblad operators [see Eq. (3)]
${L^{(1)}_{21}}$ $={e^{\frac{\delta}{2{k_{B}}T}}}{({L^{(2)}_{21}})^{\dagger}}$
$={e^{\frac{\delta}{4{k_{B}}T}}}\sqrt{\frac{\gamma}{2}\text{csch}\left(\frac{\delta}{2{k_{B}}T}\right)}|{0}\rangle\langle{1}|\penalty 10000\ ,$ (30)
where $\delta={\varepsilon_{1}}-{\varepsilon_{2}}$ is the gap of the two-level system, and $\gamma$ is the strength of the coupling between system and environment.
The spectral decomposition of the Liouvillian is given in Table 1, with left eigenmatrices ${\texttt{L}_{s}}={U_{1}}{\texttt{L}^{\prime}_{s}}{U_{1}^{\dagger}}$, and right eigenmatrices ${\texttt{R}_{s}}={U_{1}}{\texttt{R}^{\prime}_{s}}{U_{1}^{\dagger}}$, for $s\in\{1,\ldots,4\}$. The slowest decaying mode is related to the complex pair of eigenvalues $\lambda_{2}$} and ${\lambda_{3}}={\lambda^{*}_{2}}$. In particular, we verify that ${\texttt{L}_{1}}=|0\rangle\langle{0}|+|1\rangle\langle{1}|$ is the identity matrix, while ${\texttt{R}_{1}}={p_{+}}|0\rangle\langle{0}|+{p_{-}}|1\rangle\langle{1}|$ is the steady state of the dynamics, with ${p_{\mp}}=(1/2){e^{\mp\delta/(2{k_{B}}T)}}\text{sech}(\delta/(2{k_{B}}T))$. It can be proved that ${p_{-}}={e^{-{\varepsilon_{1}}/({k_{B}}T)}}/Z$, and ${p_{+}}={e^{-{\varepsilon_{2}}/({k_{B}}T)}}/Z$ are Boltzmann weights related to a thermal state at finite temperature $T$, with $Z=\text{Tr}({e^{-H/{k_{B}}T}})$ being the partition function.
Table 1: Spectral decomposition of the Liouvillian operator that governs the effective nonunitary dynamics of the two-level system described in Sec. V.1. We recall that ${\texttt{L}_{s}}={U_{1}}{\texttt{L}^{\prime}_{s}}{U_{1}^{\dagger}}$ and ${\texttt{R}_{s}}={U_{1}}{\texttt{R}^{\prime}_{s}}{U_{1}^{\dagger}}$, for $s\in\{1,\ldots,4\}$, with ${U_{1}}=i(|{1}\rangle\langle{0}|-|{0}\rangle\langle{1}|)$. Here, we introduce the Boltzmann weights ${p_{\mp}}=(1/2){e^{\mp\delta/(2{k_{B}}T)}}\text{sech}(\delta/(2{k_{B}}T))$, where $\delta={\varepsilon_{1}}-{\varepsilon_{2}}$ is the gap of the two-level system, $\gamma$ is the strength of the coupling between system and environment, $T$ is the temperature of the bosonic reservoir, and $k_{B}$ is the Boltzmann constant.
Eigenvalue Eigenmatrix
${\lambda_{1}}=0$ ${\text{L}^{\prime}_{1}}=|{0}\rangle\langle{0}|+|{1}\rangle\langle{1}|$
${\text{R}^{\prime}_{1}}={p_{-}}|{0}\rangle\langle{0}|+{p_{+}}|{1}\rangle\langle{1}|$
$\begin{aligned} &{\lambda_{2}}=\\ &-\frac{\gamma}{2}\coth\left(\frac{\delta}{2{k_{B}}T}\right)-i\delta\end{aligned}$ ${\text{L}^{\prime}_{2}}=|{0}\rangle\langle{1}|$
${\text{R}^{\prime}_{2}}=|{0}\rangle\langle{1}|$
$\begin{aligned} &{\lambda_{3}}={\lambda^{*}_{2}}=\\ &-\frac{\gamma}{2}\coth\left(\frac{\delta}{2{k_{B}}T}\right)+i\delta\end{aligned}$ ${\text{L}^{\prime}_{3}}={\text{L}^{\prime\dagger}_{2}}=|{1}\rangle\langle{0}|$
${\text{R}^{\prime}_{3}}={\text{R}^{\prime\dagger}_{2}}=|{1}\rangle\langle{0}|$
${\lambda_{4}}=-\gamma\coth\left(\frac{\delta}{2{k_{B}}T}\right)$ ${\text{L}^{\prime}_{4}}={p_{-}}|1\rangle\langle{1}|-{p_{+}}|0\rangle\langle{0}|$
${\text{R}^{\prime}_{4}}=|{1}\rangle\langle{1}|-|{0}\rangle\langle{0}|$
In this setting, since $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))\neq 0$ and $\text{Tr}({\texttt{L}_{3}}\rho({t_{0}}))\neq 0$, we apply our protocol to accelerate relaxation to equilibrium, and also induce the quantum Mpemba effect. We recall that the spectrum of the Hamiltonian is listed in descending order, ${\varepsilon_{1}}>{\varepsilon_{2}}$. This implies that the eigenvalues of the steady state are sorted in ascending order, that is, $\Sigma={U_{1}^{\dagger}}{\texttt{R}_{1}}{U_{1}}=\text{diag}({p_{-}},{p_{+}})$, with ${p_{-}}<{p_{+}}$, for all $T>0$. In this case, we verify that the spectrum of the initial state must be sorted in descending order to maximize its distance from equilibrium evaluated using the Hilbert-Schmidt distance, trace distance, and quantum relative entropy. To do so, we consider the permutation matrix ${P_{\pi}}=|{0}\rangle\langle{1}|+|{1}\rangle\langle{0}|$, such that ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}(1,0)=|0\rangle\langle{0}|$. Therefore, the transformed state becomes [see Eq. (III)]
${\rho^{\prime}}({t_{0}})={U_{1}}{P_{\pi}}{D}{P_{\pi}^{\dagger}}{U_{1}^{\dagger}}=|1\rangle\langle{1}|\penalty 10000\ .$ (31)
It can be shown that the dressed initial state in Eq. (31) leads to the suppression of the slowest decaying mode, i.e., $\text{Tr}({\texttt{L}_{2}}{\rho^{\prime}}({t_{0}}))=\langle{1}|{U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}|{1}\rangle=0$, and $\text{Tr}({\texttt{L}_{3}}{\rho^{\prime}}({t_{0}}))=\langle{1}|{U_{1}}{\texttt{L}^{\prime}_{3}}{U_{1}^{\dagger}}|{1}\rangle=0$. This ensures that the relaxation process is accelerated.
To investigate the quantum Mpemba effect, we analytically evaluate the instantaneous states $\rho(t)={e^{t\mathcal{L}}}[{\rho}({t_{0}})]$, and ${\rho^{\prime}}(t)={e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})]$. On the one hand, we find that $\rho(t)=(1/2)(\mathbb{I}+\vec{r}(t)\cdot\vec{\sigma})$, where the Bloch vector $\vec{r}(t)=({r_{x}}(t),{r_{y}}(t),{r_{z}}(t))$ has the following time-dependent components
${r_{x}}(t)$ $=-{e^{-\frac{\gamma{t}}{2}\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}\sin(\delta{t})\penalty 10000\ ,$ (32)
${r_{y}}(t)$ $=-{e^{-\frac{\gamma{t}}{2}\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}\cos(\delta{t})\penalty 10000\ ,$ (33)
${r_{z}}(t)$ $=\left(1-{e^{-\gamma{t}\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}\right)\text{tanh}\left(\frac{\delta}{2{k_{B}}T}\right)\penalty 10000\ .$ (34)
Here, $\vec{\sigma}=({\sigma_{x}},{\sigma_{y}},{\sigma_{z}})$ is the vector of Pauli matrices. On the other hand, we obtain ${\rho^{\prime}}(t)={e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})]=(1/2)(\mathbb{I}+{\vec{r}{\,{}^{\prime}}}(t)\cdot\vec{\sigma})$, with the time-dependent Bloch vector ${\vec{r}{\,{}^{\prime}}}(t)=({r^{\prime}_{x}}(t),{r^{\prime}_{y}}(t),{r^{\prime}_{z}}(t))$, where
${r^{\prime}_{x}}(t)$ $={r^{\prime}_{y}}(t)=0\penalty 10000\ ,$ (35)
${r^{\prime}_{z}}(t)$ $=\frac{2{e^{\frac{\delta}{{k_{B}}T}}}\left(1-{e^{-\gamma{t}\,\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}\right)}{{e^{\frac{\delta}{{k_{B}}T}}}+1}-1\penalty 10000\ .$ (36)
In turn, the steady state can be recast as ${\texttt{R}_{1}}=(1/2)(\mathbb{I}+{\vec{r}_{\text{ss}}}\cdot\vec{\sigma})$, with the time-independent Bloch vector ${\vec{r}_{\text{ss}}}=({r_{\text{ss}}^{x}},{r_{\text{ss}}^{y}},{r_{\text{ss}}^{z}})$, where
${r_{\text{ss}}^{x}}={r_{\text{ss}}^{y}}=0\penalty 10000\ ,\quad{r_{\text{ss}}^{z}}=\text{tanh}\left(\frac{\delta}{2{k_{B}}T}\right)\penalty 10000\ .$ (37)
Next, we address the Hilbert-Schmidt distance, trace distance, and quantum relative entropy for these states. We find that the HSD and TD are related as ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})=\sqrt{2}\,{\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$, and also ${\mathcal{D}_{\text{HSD}}}({\rho^{\prime}}(t),{\texttt{R}_{1}})=\sqrt{2}\,{\mathcal{D}_{\text{TD}}}({\rho^{\prime}}(t),{\texttt{R}_{1}})$, with
${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})=\frac{1}{2}{e^{-\frac{{\gamma}{t}}{2}\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}\times$
$\sqrt{1+{e^{-{\gamma}{t}\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}{\text{tanh}^{2}}\left(\frac{\delta}{2{k_{B}}T}\right)}\penalty 10000\ ,$ (38)
and
${\mathcal{D}_{\text{TD}}}({\rho^{\prime}}(t),{\texttt{R}_{1}})=\frac{{e^{-{\gamma}{t}\,\text{coth}\left(\frac{\delta}{2{k_{B}}T}\right)}}{\text{sech}^{2}}\left(\frac{\delta}{2{k_{B}}T}\right)}{2\left(1-\text{tanh}\left(\frac{\delta}{2{k_{B}}T}\right)\right)}\penalty 10000\ .$ (39)
The analytical expressions for the relative entropies $S(\rho(t),{\texttt{R}_{1}})$ and $S({\rho^{\prime}}(t),{\texttt{R}_{1}})$ are too long to be reported here. We note that QRE is evaluated with the help of the following expression de Sousa and Pires
$S(\varrho,{\texttt{R}_{1}})=|\vec{u}\,|\,\text{ln}\left(\sqrt{\frac{1+|\vec{u}\,|}{1-|\vec{u}\,|}}\,\right)$
$+\text{ln}\left(\sqrt{\frac{1-{|\vec{u}\,|^{2}}}{1-{|{\vec{r}_{\text{ss}}}|^{2}}}}\,\right)-\text{ln}\left(\sqrt{\frac{1+|{\vec{r}_{\text{ss}}}|}{1-|{\vec{r}_{\text{ss}}}|}}\right)\frac{(\vec{u}\cdot{\vec{r}_{\text{ss}}})}{|{\vec{r}_{\text{ss}}}|}\penalty 10000\ ,$ (40)
which holds for any single-qubit state $\varrho=(1/2)(\mathbb{I}+\vec{u}\cdot\vec{\sigma})$. We recall that the Bloch vectors $\vec{r}(t)$, ${\vec{r}{\,{}^{\prime}}}(t)$, and $\vec{r}_{\text{ss}}$} were described in Eqs. (32)–(37).
In Fig. 2, we show the plots for the trace distance [see Fig. 2(a)], and quantum relative entropy [see Fig. 2(b)] related to the states $\rho(t)$, ${\rho^{\prime}}(t)$, and $\texttt{R}_{1}$, as a function of the dimensionless parameter $\gamma{t}$. We set the parameters $\delta={\varepsilon_{1}}-{\varepsilon_{2}}=1$, $T=1$, and ${k_{B}}=1$. The plots clearly show that the relaxation to the equilibrium is accelerated as a consequence of using the transformed state ${\rho^{\prime}}({t_{0}})$. The plots show the occurrence of the quantum Mpemba effect, which is testified by the crossover between the relaxation curves for the selected distinguishability measures.
Figure 2: (Color online) The quantum Mpemba effect in the dynamics of the two-level system described in Sec. V.1. Here we use $\delta={\varepsilon_{1}}-{\varepsilon_{2}}=1$, $T=1$, and ${k_{B}}=1$. To investigate the relaxation of the instantaneous states $\rho(t)$ and ${\rho^{\prime}}(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) trace distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $\gamma{t}$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
Figure 3: (Color online) The quantum Mpemba effect in the dynamics of the transverse field Ising (TFI) model with open boundary conditions [see details in Sec. V.2], where $N=5$ spins, $h=J/2$, $\gamma=1$, and $T=0.1$. To investigate the relaxation of the instantaneous states $\rho(t)$ and ${\rho^{\prime}}(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) Hilbert-Schmidt distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$; (c) trace distance, ${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $Jt$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
V.2 Many-body quantum systems